# Mingkai Liao

### v1

@@ 5.3 Synthetic Benchmark: Performance across Graph Sizes
We further evaluate PACIFIER on a controlled synthetic benchmark to examine generalization across graph sizes and to provide an apples-to-apples comparison with competing methods under identical instance distributions. We generate two-echo-chamber graphs following the procedure in Sec. 4.2, and report results on six node-size ranges: $\{30$50, 5050,\,50100, 100100,\,100200, 200200,\,200300, 300300,\,300400, 400400,\,400500}500\}. For each size range, we sample $100$ graphs, resulting in $600$ test instances in total. We compare PACIFIER against all baselines on four tasks: MI, MI-cost, ME, and ME-cost. In each bar chart, the x-axis corresponds to the node-size range and each bar reports the average score over the $100$ graphs in that range (with error bars showing the standard deviation when available).
@@
(b) MI-cost on synthetic graphs.
(c) ME on synthetic graphs.
(d) ME-cost on synthetic graphs.
Figure 2: Synthetic benchmark results across six node-size ranges (30–50, 50–100, 100–200, 200–300, 300–400, 400–500), with $100$ graphs per range. Bars compare PACIFIER with all baselines on four tasks: MI, MI-cost, ME, and ME-cost. Lower is better if the reported metric is polarization/ANP; higher is better if it is a score (%). (See the main text for the exact metric definition used in these plots.)
@@ Discussion.
As shown in Fig. 2, the synthetic benchmark reveals three qualitatively different regimes across tasks, which together clarify when long-horizon reinforcement learning is essential and when myopic learning already suffices.
@@
(1) MI (unweighted): PACIFIER matches the linear-model optimum. From Fig. 2(a), we observe that on the canonical MI task, both PACIFIER-RL and PACIFIER-Greedy achieve performance nearly indistinguishable from BOMP, the classical model-aware algorithm that explicitly exploits the linear structure of the FJ steady-state solution. Across all graph-size ranges (30–50 up to 400–500), the two PACIFIER variants closely track the BOMP curve and significantly outperform purely structural heuristics such as PageRank, ExtremeExpressed, ExtremeNeighbours, and random selection. This indicates that the learned representation successfully captures the same influence structure that BOMP leverages analytically, despite not using closed-form matrix inversions during deployment. Moreover, the negligible gap between PACIFIER-RL and PACIFIER-Greedy suggests that in this linear, unweighted setting, the immediate polarization-reduction signal is already highly aligned with the long-horizon objective; thus, myopic value estimation is sufficient to approximate near-optimal action ranking.
(2) MI-cost: RL and Greedy remain close, but both dominate heuristics. As illustrated in Fig. 2(b), when heterogeneous intervention costs are introduced (MI-cost), both PACIFIER variants continue to substantially outperform classical heuristics. Although BOMP is no longer directly optimal under cost-weighting, PACIFIER-RL and PACIFIER-Greedy remain very close to each other across all scales, indicating that even with cost-aware trade-offs, the immediate reward remains a strong proxy for long-term utility. Importantly, both variants consistently outperform structural baselines, demonstrating that the encoder–decoder architecture internalizes cost-sensitive influence patterns without relying on linear closed-form solutions.
(3) ME and ME-cost: long-horizon RL becomes essential. In contrast, Fig. 2(c) and Fig. 2(d) show that under ME and especially ME-cost, PACIFIER-RL decisively outperforms PACIFIER-Greedy across all graph sizes. The gap between RL and Greedy is persistent and substantial, while both PACIFIER variants clearly dominate all traditional baselines. This indicates that fixing expressed opinions induces stronger sequential and interaction effects: the benefit of intervening on one node depends critically on how the system re-equilibrates and how future feasible actions are reshaped by previously fixed constraints. Such dependencies cannot be captured by immediate reward regression alone, but are naturally learned through bootstrapped Q-learning. The stability of this RL advantage across graph scales confirms that long-horizon credit assignment is fundamentally important for ME-style interventions rather than being a small-graph artifact.
@@
(4) Scalability and regime-dependent optimality. Across all four subfigures in Fig. 2, performance remains remarkably stable as the graph size increases by an order of magnitude. The relative ordering between methods does not change with scale, demonstrating strong inductive generalization from small training graphs to larger synthetic instances. Collectively, the results illustrate a regime-dependent phenomenon: when the task admits a strong linear analytical structure (as in MI), PACIFIER recovers near-optimal behavior comparable to BOMP; when that structure is weakened by cost heterogeneity or by expressed-opinion constraints (MI-cost, ME, ME-cost), learning-based long-horizon reasoning provides clear and consistent gains.
5.4 Real-world Datasets (origin)
@@
We report results on the real-world polarized topic graphs described in Sec. 5.1. For each task, we provide: (i) a per-dataset bar chart summarizing the ANP/AUC-style metric, (ii) a heatmap overview across datasets and methods, and (iii) aggregated polarization trajectories.
@@
Figure 4: Real datasets (MI-cost): per-dataset bar comparison (lower is better).
@@
Figure 6: Real datasets (ME-cost): per-dataset bar comparison (lower is better).
@@
(b) MI-cost.
(c) ME.
(d) ME-cost.
Figure 7: Real datasets (origin): heatmap summaries for MI/MI-cost/ME/ME-cost. Lower is better.
@@
To qualitatively illustrate when polarization reduction happens along the intervention horizon, we visualize polarization trajectories on three representative datasets for each task. Specifically, we select (i) the dataset with the smallest number of nodes, (ii) the largest dataset whose size is below a threshold (e.g., n<20,000n<20{,}000), and (iii) the dataset with the largest number of nodes. This selection covers small/medium/large regimes while keeping the main text concise. The complete set of trajectory plots for all datasets is provided in Appendix.
Figure 8: Real datasets (MI): polarization trajectories on three representative datasets (small / medium / large by node count). Lower and earlier is better. Full trajectories for all datasets are in Appendix.
Figure 9: Real datasets (MI-cost): polarization trajectories on three representative datasets (small / medium / large by node count). Lower and earlier is better. Full trajectories for all datasets are in Appendix.
Figure 10: Real datasets (ME): polarization trajectories on three representative datasets (small / medium / large by node count). Lower and earlier is better. Full trajectories for all datasets are in Appendix.
Figure 11: Real datasets (ME-cost): polarization trajectories on three representative datasets (small / medium / large by node count). Lower and earlier is better. Full trajectories for all datasets are in Appendix.
@@ 5.4.1 Discussion
We next analyze the real-world results in detail. For each intervention setting (MI, MI-cost, ME, and ME-cost), we discuss the empirical findings from three complementary perspectives: (i) per-dataset bar comparisons, (ii) heatmap-level performance patterns across datasets, and (iii) representative polarization trajectories that reveal the dynamic evolution of polarization. We examine each task individually and then summarize the broader trends observed across settings.
@@
On the MI setting, PACIFIER does not outperform the strongest non-PACIFIER baselines on the real-world graphs. As shown in Fig. 3, BOMP achieves the lowest accumulated metric on the majority of datasets, with an average AUC of 0.258 compared to 0.298 for PACIFIER-RL and 0.305 for PACIFIER-Greedy. Across all 15 datasets, PACIFIER-RL does not achieve the best result on any instance, and the average relative gap to the strongest baseline is approximately 8%, with larger discrepancies observed on some small-scale graphs.
The heatmap in Fig. 7(a) further confirms this pattern. The advantage of classical structural heuristics such as BOMP and ExtremeNeighbours is consistent across small, medium, and large graphs. Although the performance gap slightly narrows as the graph size increases (e.g., on large retweet networks), PACIFIER remains systematically behind the best-performing baseline in all size regimes.
The representative trajectory plots in Fig. 8 provide additional insight into this behavior. On small graphs (e.g., follow_germanwings), PACIFIER reduces polarization more slowly in the early intervention stages and fails to catch up later, leading to a consistently higher trajectory. On medium and large graphs (e.g., retweet_onedirection and retweet_mothersday), PACIFIER follows a trajectory that remains above the strongest baseline throughout the horizon, indicating that the gap is not due to late-stage instability but rather a persistent structural disadvantage. Overall, under the MI objective on real-world graphs, classical heuristics appear better aligned with the polarization structure than the learned PACIFIER strategy.
MI-cost.
The picture changes dramatically under the cost-aware MI setting. As shown in Fig. 4, both PACIFIER-RL and PACIFIER-Greedy consistently achieve the lowest accumulated metric across all 15 datasets. PACIFIER-RL improves over the strongest non-PACIFIER baseline on every dataset, with a mean relative gain of approximately 37% (median 41%), while PACIFIER-Greedy achieves a comparable mean improvement of about 38%. This represents a complete reversal of the trend observed in the cost-free MI setting.
The heatmap in Fig. 7(b) further highlights the stability of this advantage. Unlike the MI case, where classical heuristics dominate, PACIFIER uniformly outperforms all baselines across small, medium, and large graphs. Moreover, the relative improvement increases with graph size: the average gain is around 27% on small graphs, rises to nearly 40% on medium graphs, and exceeds 42% on large-scale networks. This clear scalability pattern suggests that PACIFIER is particularly effective when intervention decisions must be made under budget constraints in large real-world graphs.
The representative trajectory plots in Fig. 9 provide insight into the source of this improvement. On small graphs (e.g., follow_germanwings), PACIFIER reduces polarization more aggressively from the early stages, resulting in a visibly lower trajectory throughout the intervention horizon. On medium and large graphs (e.g., retweet_onedirection and retweet_mothersday), the gap becomes even more pronounced: PACIFIER maintains a substantially lower trajectory at nearly every step, rather than merely converging to a better final value. This indicates that the advantage stems from consistently better budget allocation decisions, leading to sustained polarization suppression over time.
@@
Under the ME objective, PACIFIER-RL consistently achieves the best performance across all real-world datasets. As illustrated in Fig. 5, PACIFIER-RL attains the lowest accumulated metric on all 15 datasets, with an average AUC of 0.105, outperforming the strongest non-PACIFIER baseline (typically PageRank) by approximately 15% on average. The improvement is particularly pronounced on certain small graphs (up to 31% relative gain), while even on the most challenging instances the advantage remains positive.
The heatmap in Fig. 7(c) confirms that this dominance is stable across scales. PACIFIER-RL achieves a 100% win rate in small-, medium-, and large-scale regimes, with mean relative gains of roughly 19%, 14%, and 13%, respectively. Although the improvement magnitude slightly decreases as graph size grows, the method remains consistently superior across all size buckets. In contrast, PACIFIER-Greedy exhibits noticeably less stability, winning only 60% of datasets overall and showing performance degradation on several small and medium graphs, highlighting the benefit of the reinforcement learning component.
The representative trajectory plots in Fig. 10 further clarify the source of the improvement. On small graphs (e.g., follow_germanwings), PACIFIER-RL rapidly suppresses polarization in early stages, creating a clear separation from structural heuristics. On medium graphs (e.g., retweet_onedirection), the advantage is maintained steadily across the entire horizon. Even on very large graphs (e.g., retweet_mothersday), PACIFIER-RL preserves a consistently lower trajectory than PageRank, whereas PACIFIER-Greedy occasionally loses its advantage. Overall, under the ME objective, the learned RL policy demonstrates strong and robust generalization on real-world networks.
ME-cost.
Under the cost-aware ME objective, PACIFIER-RL achieves the strongest overall performance among all methods. As shown in Fig. 6, PACIFIER-RL attains the lowest accumulated metric on all 15 datasets, with an average AUC of 0.075, substantially outperforming the strongest non-PACIFIER baseline by approximately 39% on average. The relative gain remains positive on every dataset, with improvements ranging from about 29% to nearly 57%, indicating highly consistent dominance under budget constraints.
The heatmap in Fig. 7(d) further demonstrates the robustness of this advantage. PACIFIER-RL achieves a 100% win rate across small-, medium-, and large-scale graphs, with mean improvements of roughly 45% on small graphs, 36% on medium graphs, and 38% on large graphs. Unlike the cost-free ME setting, where the improvement magnitude slightly decreases with scale, the cost-aware formulation preserves a strong and stable advantage across all regimes. In contrast, PACIFIER-Greedy exhibits considerable instability, achieving only partial wins (approximately 47% overall) and even negative improvements on several medium and large datasets, highlighting the importance of the learned RL policy in budget-sensitive scenarios.
The representative trajectory plots in Fig. 11 illustrate the dynamic source of this superiority. On small graphs (e.g., follow_germanwings), PACIFIER-RL dramatically lowers polarization from the earliest intervention stages, creating a large and persistent separation from structural heuristics. On medium and large graphs (e.g., retweet_onedirection and retweet_mothersday), the RL policy maintains a consistently lower trajectory throughout the horizon, rather than merely achieving a better final value. This sustained gap suggests that PACIFIER-RL effectively prioritizes influential nodes under budget constraints, resulting in systematically improved polarization mitigation over time.
5.5 Real Datasets: Extended Settings
[…]
AUC summary (bar; by dataset).
The bar plots summarize each method’s performance on every real dataset under the same budget, where lower is better. Since the plots are wide, we place one bar chart per line.

### v2

@@ 5.4 Synthetic Benchmark: Performance across Graph Sizes
We further evaluate PACIFIER on a controlled synthetic benchmark to examine generalization across graph sizes and to provide an apples-to-apples comparison with competing methods under identical instance distributions. We generate two-echo-chamber graphs following the procedure in Sec. 4.1, and report results on six node-size ranges: $\{30$50, 5050,\,50100, 100100,\,100200, 200200,\,200300, 300300,\,300400, 400400,\,400500}500\}. For each size range, we sample $100$ graphs, resulting in $600$ test instances in total. We compare PACIFIER against all baselines on three tasks: MI, ME, and node-removal. In each bar chart, the x-axis denotes the node-size range, and each bar reports the average score over the $100$ graphs in that range, with error bars indicating the standard deviation when available.
@@
(b) ME on synthetic graphs.
(c) node-removal on synthetic graphs.
Figure 2: Synthetic benchmark results across six node-size ranges (30–50, 50–100, 100–200, 200–300, 300–400, 400–500), with $100$ graphs per range. Bars compare PACIFIER with all baselines on three tasks: MI, ME, and node-removal. The score represented by each bar height is the ANP score reported in this paper, and lower values indicate better performance.
@@ Discussion.
As shown in Fig. 2, the synthetic benchmark reveals three distinct behavioral regimes across tasks, which together clarify when simple myopic reasoning is sufficient and when long-horizon reinforcement learning becomes essential.
@@
(1) MI: PACIFIER matches the linear-model optimum. From Fig. 2(a), we observe that on the canonical MI task, both PACIFIER-RL and PACIFIER-Greedy achieve performance very close to BOMP, the classical model-aware algorithm that explicitly exploits the linear structure of the FJ steady-state solution. Across all graph-size ranges from 30–50 up to 400–500, the two PACIFIER variants closely track the BOMP curve and consistently outperform structural heuristics such as PageRank, ExtremeExpressed, ExtremeNeighbours, and random selection. This indicates that the learned representation successfully captures the same influence structure that BOMP leverages analytically, despite not relying on closed-form matrix operations at deployment time. Moreover, the negligible gap between PACIFIER-RL and PACIFIER-Greedy suggests that in this linear and comparatively well-structured setting, the immediate polarization-reduction signal is already well aligned with the long-term objective. As a result, myopic value estimation is sufficient to recover near-optimal action rankings.
(2) ME: RL shows stronger scale generalization than Greedy on synthetic graphs. In contrast, Fig. 2(b) shows a markedly different pattern for ME. Here, PACIFIER-RL consistently achieves the best performance across all graph sizes, while PACIFIER-Greedy is noticeably weaker and remains much closer to the stronger classical baselines. Relative to Greedy, RL exhibits clearly stronger scale generalization on the synthetic benchmark, that is, under matched synthetic training and test distributions without considering the training-distribution gap that may arise on real datasets. The gap between RL and Greedy is persistent rather than incidental, indicating that once interventions operate on expressed opinions, the action value can no longer be inferred reliably from immediate single-step gains alone. Fixing a node’s expressed opinion changes not only the current equilibrium, but also the downstream feasibility and utility of future interventions, creating strong sequential dependencies. Such effects are difficult to capture with purely myopic regression, but are naturally handled by bootstrapped Q-learning. The stability of this RL-over-Greedy advantage across all six graph-size ranges therefore suggests that long-horizon credit assignment is important for ME-style interventions rather than being only a small-graph artifact.
(3) Node-removal: RL clearly outperforms Greedy. Fig. 2(c) presents an even clearer regime. For node-removal, PACIFIER-RL achieves the lowest score in every size range, with a consistent margin over all baselines. At the same time, PACIFIER-Greedy remains the second-best method overall and also substantially outperforms BOMP, PageRank, ExtremeExpressed, ExtremeNeighbours, and random selection. More importantly, RL is clearly better than Greedy in this setting. Compared with MI, this larger RL–Greedy gap suggests that node-removal introduces stronger combinatorial and long-range intervention effects: removing one node can fundamentally alter the subsequent graph topology, influence pathways, and the marginal utility of later actions. Therefore, while myopic learning already captures useful local signals, full reinforcement learning is much better suited to reasoning over the nonlinear sequential consequences induced by structural interventions.
@@
(4) Scalability and regime-dependent optimality. Across all three subfigures in Fig. 2, the relative ordering between methods remains highly stable as the graph size increases from 30–50 nodes to 400–500 nodes. This demonstrates strong inductive generalization from smaller training graphs to substantially larger synthetic instances. Taken together, the results reveal a clear regime-dependent picture: when the task admits a strong linear analytical structure as in MI, PACIFIER recovers near-optimal behavior comparable to BOMP. When interventions induce stronger sequential or structural dependencies, as in ME and especially node-removal, long-horizon reinforcement learning provides clear and consistent gains.
5.5 Real-world Datasets
@@
We report results on the real-world polarized topic graphs described in Sec. 5.1. We focus on three representative intervention settings: the two canonical topology-preserving tasks, MI and ME, and the topology-altering setting node_removal. For each task, we provide: (i) a per-dataset bar chart summarizing the ANP/AUC-style metric, (ii) a heatmap overview across datasets and methods, and (iii) representative polarization trajectories illustrating when polarization reduction occurs along the intervention horizon.
@@
Figure 5: Real datasets (node_removal): per-dataset bar comparison (lower is better).
@@
(b) ME.
(c) node_removal.
Figure 6: Real datasets: heatmap summaries for MI, ME, and node_removal. Lower is better.
@@
To qualitatively illustrate when polarization reduction happens along the intervention horizon, we visualize polarization trajectories on three representative datasets for each task. Specifically, we select (i) the dataset with the smallest number of nodes, (ii) the largest dataset whose size is below a threshold (e.g., n<20,000n<20{,}000), and (iii) the dataset with the largest number of nodes. This selection covers small/medium/large regimes while keeping the main text concise. The complete set of trajectory plots for all datasets is provided in the Appendix.
Figure 7: Real datasets (MI): polarization trajectories on three representative datasets (small / medium / large by node count). Lower and earlier is better. Full trajectories for all datasets are in the Appendix.
Figure 8: Real datasets (ME): polarization trajectories on three representative datasets (small / medium / large by node count). Lower and earlier is better. Full trajectories for all datasets are in the Appendix.
Figure 9: Real datasets (node_removal): polarization trajectories on three representative datasets (small / medium / large by node count). Lower and earlier is better. Full trajectories for all datasets are in the Appendix.
@@ 5.5.1 Discussion
We next analyze the real-world results in detail. For each intervention setting (MI, ME, and node_removal), we discuss the empirical findings from three complementary perspectives: (i) per-dataset bar comparisons, (ii) heatmap-level performance patterns across datasets, and (iii) representative polarization trajectories that reveal the temporal evolution of polarization. Together, these three tasks allow us to compare PACIFIER across both topology-preserving and topology-altering regimes.
Evaluation protocol note.
Across all three real-world tasks (MI, ME, and node_removal), we evaluate on the same set of 15 polarized topic graphs. However, BOMP is only available on 10 of these 15 datasets, because its computation becomes prohibitively expensive on the five largest graphs. As a result, any average reported for BOMP is computed on this 10-dataset subset only, and is therefore not directly comparable to averages of methods such as PACIFIER that are reported over all 15 datasets. Accordingly, in the task-specific discussion below, we distinguish subset-based average comparisons from full-coverage results, and we place primary emphasis on per-dataset comparisons against the strongest available non-PACIFIER baseline.
@@
On the real-world MI benchmark, PACIFIER does not surpass BOMP when the latter is computationally feasible, which is expected since MI is precisely the regime where analytical structure is the cleanest and BOMP can be regarded as the strongest classical solver. In this sense, the real-world MI results should not be interpreted as a failure of the framework, but rather as a test of how much of this analytical decision structure PACIFIER can recover from synthetic training and transfer to heterogeneous real graphs.
As noted above, BOMP is only available on 10 out of the 15 real-world datasets, because its computation becomes prohibitively expensive on the five largest graphs. On this 10-dataset BOMP-comparable subset, BOMP remains the strongest method on average, achieving an average AUC of 0.258. PACIFIER-Greedy and PACIFIER-RL obtain average AUCs of 0.281 and 0.282, respectively, and neither variant outperforms BOMP on any of these 10 datasets. This confirms that, in the canonical linear MI regime, the strongest analytical solver still retains a clear advantage whenever it is available.
At the same time, these results also suggest a mild but non-negligible synthetic-to-real gap. Although our synthetic graph generator explicitly uses two echo chambers to better align with the opposing-camp structure observed in polarized real networks, this approximation is still not sufficient to fully match the heterogeneity of real MI instances. As a result, PACIFIER can recover a substantial portion of the useful decision structure learned in synthetic environments, but transferring that behavior to real-world graphs remains somewhat imperfect.
However, since BOMP is already effectively the strongest classical reference for MI, the more practically relevant comparison is between PACIFIER and the remaining scalable non-PACIFIER baselines. Under this excluding-BOMP view, the picture becomes substantially more favorable to PACIFIER. On the same 10-dataset subset, PACIFIER-Greedy is the best method on average among all non-BOMP methods (0.281), with PACIFIER-RL a close second (0.282), both outperforming Random (0.284) and the other heuristic baselines on average. At the dataset level, PACIFIER-RL outperforms the strongest excluding-BOMP baseline on 5 out of 10 datasets, while PACIFIER-Greedy does so on 3 out of 10, with mean relative gaps of only 2.51% and 2.49%, respectively. This indicates that although PACIFIER does not match BOMP, it already becomes highly competitive once the comparison is restricted to non-analytical baselines.
The same pattern also appears under the full-coverage evaluation that excludes BOMP and uses all 15 datasets. In this setting, Random is the strongest baseline on average with an AUC of 0.295, while PACIFIER-Greedy and PACIFIER-RL achieve 0.297 and 0.298, respectively. Relative to the strongest available excluding-BOMP baseline on each dataset, PACIFIER-RL wins on 6 out of 15 datasets and PACIFIER-Greedy wins on 4 out of 15, with mean relative gaps of 3.21% and 3.37%, respectively. Thus, even though a synthetic-to-real transfer gap remains, the gap is moderate rather than severe when PACIFIER is compared against the scalable baselines that are actually available across the full benchmark.
Several representative datasets further support this interpretation. On follow_germanwings, PACIFIER-RL ranks second overall behind BOMP (0.129 vs. 0.125), and it also outperforms the strongest excluding-BOMP baseline, ExtremeExpressed (0.129 vs. 0.131). On retweet_onedirection, PACIFIER-RL is again second only to BOMP (0.239 vs. 0.226), while outperforming the strongest excluding-BOMP baseline, ExtremeNeighbours (0.239 vs. 0.253). At the same time, difficult cases such as retweet_mothersday show that this transfer remains non-trivial on highly heterogeneous large-scale graphs.
Overall, the MI results are best interpreted as evidence of a meaningful but bounded transfer gap. BOMP remains the strongest method whenever it is feasible, which is consistent with MI being the most analytically structured setting. But once the focus shifts to the more practical comparison against scalable non-PACIFIER baselines, PACIFIER becomes highly competitive: it is among the best methods on average, wins on a substantial fraction of datasets, and clearly preserves a significant portion of the useful decision structure learned from synthetic training.
@@
The real-world results under the ME objective are substantially stronger than those under MI. Rather than merely remaining competitive, PACIFIER consistently outperforms all classical non-PACIFIER baselines across the benchmark. This suggests that ME is a regime in which the intervention policy learned from synthetic graphs transfers to real networks particularly well, and where classical analytical or structural heuristics no longer define the performance frontier.
This conclusion already holds on the 10-dataset subset where BOMP is computationally feasible. On this subset, PACIFIER-Greedy and PACIFIER-RL achieve average AUCs of 0.0949 and 0.0956, respectively, clearly outperforming the strongest non-PACIFIER baseline, PageRank, at 0.1205, and also substantially outperforming BOMP at 0.2073. At the dataset level, both PACIFIER variants outperform the strongest non-PACIFIER baseline on all 10 datasets: PACIFIER-Greedy achieves a 10/10 win rate with a mean relative improvement of 21.55%, and PACIFIER-RL also achieves a 10/10 win rate with a mean relative improvement of 20.96%. Therefore, unlike MI, ME does not appear to be a setting where analytical baselines retain a decisive real-world advantage.
At the same time, the relationship between PACIFIER-RL and PACIFIER-Greedy is also informative. On fully aligned synthetic benchmarks, where the training and test distributions match closely, RL exhibits stronger scale generalization than Greedy, indicating that long-horizon reasoning is indeed beneficial for ME-style interventions. However, once the evaluation moves to real graphs, a mild synthetic-to-real gap still remains. Even though the synthetic generator is designed to better align with polarized real networks, real ME instances still contain additional heterogeneity that is not perfectly captured by the synthetic distribution. This gap is not large enough to undermine transfer, but it does appear to compress the difference between the two PACIFIER variants: RL and Greedy remain very close on real data, with Greedy even slightly outperforming RL on the current benchmark.
The same pattern becomes even clearer under the full-coverage comparison that excludes BOMP and uses all 15 datasets. In this setting, PACIFIER-Greedy and PACIFIER-RL obtain average AUCs of 0.0978 and 0.0986, respectively, while PageRank, the strongest non-PACIFIER baseline on average, reaches 0.1227. More importantly, both PACIFIER variants outperform the strongest available non-PACIFIER baseline on every dataset in the benchmark. PACIFIER-Greedy achieves a 15/15 win rate with a mean relative improvement of 20.52%, and PACIFIER-RL also achieves a 15/15 win rate with a mean relative improvement of 19.90%. This dominance is also highly stable across graph sizes: in the small, medium, and large buckets, both variants maintain a 100% win rate, with mean gains remaining consistently around 18%–23%.
Representative datasets further illustrate this robustness. On follow_germanwings, both PACIFIER variants dramatically outperform the strongest baseline, ExtremeNeighbours, with PACIFIER-RL achieving 0.0613 and PACIFIER-Greedy 0.0619 versus 0.0911. On retweet_onedirection, both again show a large advantage over PageRank, with PACIFIER-Greedy and PACIFIER-RL obtaining 0.0864 and 0.0876, compared with 0.1228. Even on the very large retweet_mothersday graph, where the task is more challenging and the synthetic-to-real mismatch is likely more pronounced, PACIFIER-Greedy and PACIFIER-RL still clearly outperform PageRank (0.1159 and 0.1162 versus 0.1299). These examples indicate that the learned intervention strategy transfers robustly not only on small graphs, but also on large and heterogeneous real networks.
Overall, ME provides some of the clearest evidence of successful synthetic-to-real transfer in our framework. In fully aligned synthetic settings, RL shows stronger scale generalization than Greedy, confirming the value of long-horizon reasoning for this objective. On real-world graphs, a mild transfer gap remains, but it does not change the qualitative conclusion: both PACIFIER variants remain highly effective, perform very similarly to each other, and consistently dominate all non-PACIFIER baselines across the benchmark.
Node removal.
The real-world results on the node_removal task provide strong evidence that explicit sequential decision-making is necessary for effective intervention. Unlike ME, where PACIFIER-RL and PACIFIER-Greedy remain very close on real graphs, here PACIFIER-RL clearly separates itself not only from the classical baselines, but also from PACIFIER-Greedy. This suggests that node_removal is inherently more sensitive to long-horizon credit assignment: removing a node changes the future graph structure directly, and the quality of later decisions therefore depends strongly on the earlier intervention sequence.
This conclusion already holds on the 10-dataset subset where BOMP is computationally feasible. On this subset, PACIFIER-RL is the best method on average, achieving an average AUC of 0.0346, compared with 0.0483 for PageRank, 0.0636 for PACIFIER-Greedy, and 0.3261 for BOMP. At the dataset level, PACIFIER-RL outperforms the strongest available non-PACIFIER baseline on 9 out of 10 datasets, with a mean relative improvement of 18.02%. By contrast, PACIFIER-Greedy wins on only 5 out of 10 datasets, and its mean relative improvement is strongly negative overall due to several severe failure cases. Therefore, unlike MI, BOMP is not a competitive reference in this setting, and unlike ME, the gap between RL and Greedy remains clearly visible even on real-world graphs.
The same pattern becomes even clearer under the full-coverage comparison that excludes BOMP and uses all 15 datasets. In this setting, PACIFIER-RL again ranks first on average with an AUC of 0.0270, clearly outperforming PageRank, the strongest non-PACIFIER baseline, at 0.0370. More importantly, PACIFIER-RL outperforms the strongest available non-PACIFIER baseline on 14 out of 15 datasets, with a mean relative improvement of 19.16%. In contrast, PACIFIER-Greedy achieves an average AUC of 0.0534, wins on only 6 out of 15 datasets, and exhibits a strongly negative mean relative improvement overall because of several pronounced failure cases. This shows that the advantage of PACIFIER in node_removal is not merely due to using a learned scoring function, but specifically due to the RL policy’s ability to reason over the downstream structural consequences of each intervention step.
The bucketed analysis further supports this interpretation. PACIFIER-RL remains strong across all graph sizes: it wins on 3 out of 4 small graphs, 5 out of 5 medium graphs, and 6 out of 6 large graphs, with mean relative improvements of 16.80%, 20.42%, and 19.68%, respectively. Thus, its advantage is not confined to a particular scale, but persists from small real networks to the largest graphs in the benchmark. By contrast, PACIFIER-Greedy becomes increasingly unstable as graph size grows: while it remains partially competitive on small and medium graphs, it wins on only 1 out of 6 large graphs and suffers especially large degradations on several medium- and large-scale instances. This contrast strongly suggests that scale generalization in node_removal depends crucially on true sequential planning rather than myopic ranking quality alone.
Representative datasets illustrate the same pattern. On follow_germanwings, PACIFIER-RL achieves the best overall result, substantially outperforming both BOMP and PageRank (0.0811 vs. 0.1410 and 0.1548), while PACIFIER-Greedy remains competitive but clearly weaker at 0.1208. On retweet_onedirection, PACIFIER-RL again ranks first and improves over PageRank by 15.98% (0.0143 vs. 0.0170), whereas PACIFIER-Greedy deteriorates sharply to 0.1006. Even on the extremely large retweet_mothersday graph, PACIFIER-RL still achieves the best result at 0.0119, outperforming PageRank at 0.0155, while PACIFIER-Greedy remains competitive but still trails RL at 0.0139. These cases show that the RL advantage is both substantial and robust, although Greedy can still recover useful behavior on some large instances.
Overall, the node_removal results provide strong evidence that PACIFIER-RL has learned a genuinely transferable sequential intervention policy rather than a simple static heuristic. The synthetic-to-real gap does not disappear completely, but it is clearly small enough for the RL policy to transfer very effectively across heterogeneous real graphs. At the same time, the persistent gap between RL and Greedy indicates that for node_removal, unlike ME, real-world performance depends critically on modeling the long-term interaction between current interventions and future graph evolution.
5.6 Real-world Datasets in Extended Settings
We further report results on the same real-world polarized topic graphs under two extended settings: continuous and cost. Each setting contains two topology-preserving intervention tasks, namely MI and ME. To keep the presentation concise and focused, we report only the per-dataset bar charts for these extended settings. These figures summarize the trajectory-aware performance across datasets using the same ANP/AUC-style metric as in the main experiments, where lower values indicate better performance.
5.6.1 Continuous Setting
Per-dataset bar charts.
The following bar charts compare PACIFIER with all baselines on each dataset under the continuous-opinion setting. Lower values indicate better trajectory-aware performance (area-under-curve / accumulated metric).